OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4

OpenStudy (zzr0ck3r):

can you solve for y?

OpenStudy (zzr0ck3r):

put it in the form y=mx+b then m is slope and b is the y intercept

OpenStudy (anonymous):

x - 4y = 4 Get the expression into SLOPE INTERCEPT FORM: \(y = mx + b \) x - 4y = 4 subtract the x x - 4y = 4 -x -x ---- ----- -4y = -x + 4 Divide by -4 on both sides -4y = -x + 4 ---- ------ -4 -4 y = try it :)